TECHNICAL OPERATIONS MEETING #00-09 1 May 2000 Present: R. Atapattu, G. Barrick, R. Calder, W. Cruise, G. Fahlman, T. Forveille, B. Grundseth, O. Lai, E. Magnier, N. Manset, F. Menard, D. Salmon, J. Thomas, C. Veillet 1.0 OBSERVING REPORTS 1.1 Instrument: CFH12K Date: 25-27 April 2000 (Report by E. Magnier) Observers: Michael West Program: 2000AH20 Success: When the weather and seeing cooperated, they were able to obtain useful data. The first two nights were almost completely wiped out (1 hour useful). The third night has excellent weather. Problems: Poor weather for much of the run. No significant operational problems. 1.2 Instrument: CFH12K Date: 28-30 April 2000 (Report by E. Magnier) Observers: Henry McCracken Program: 2000AF52 Success: Two of the three nights had good weather with decent seeing. The observers were able to obtain good data those nights. Problems: Poor weather ruined most of the first night of the run. 2.0 PROGRESS REPORTS 2.1 SAFETY OFFICER (R. Atapattu) No report. 2.2 MECHANICAL (R. Atapattu) 2.2.1 Daycrew (Taroma, Beck, Uchima, and Wood) Unload truck and put away supplies and freight. Building checks, switch telescope air drying to 1st floor unit. Checked aluminizing for additional clean up. Placed Ln2 into elevator; cleared hatchway and loaded trash to bring to Waimea. Changed out desiccant on 1st floor. Building checks. Ordered parts for glycol pump. disconnect and connect glycol pump on mirror cooling. Assisted contractors with glycol piping. Secure cover on electrical panel. Attended safety conference on Oahu. 2.2.2 Machine Shop - Dan Sabin No report. 2.2.3 Mechanical Design - Wiley Knight No report. 2.2.4 Engineering - Rohendra Atapattu No report. 2.3 DETECTOR, ELECTRONICS & TCS GROUP (R. Calder) 2.3.1 CFHT-IR: Analyzed Photon Transfer Curve data multiple times (Greg fit linearity of multiple data sets). (Jamey, Barry, Greg) Dewar prep for shipment. (Barry, Jamey, Bob) Dewar shipped to UdeM for optics integration. (Jamey, Barry, Bob) 2.3.2 CAFE: Attend CAFE planning meeting. (Grant) 2.3.3 MegaPrime/MegaCam: Reviewed progress on MegaPrime cooling and electronics systems. (Will) Reviewed documents for and attended MegaPrime Control Software pre-pdr. (Will) Reviewed documents for and attended MegaPrime Guider CCD software/hardware FDR. (Will) Prepared re-design documents in support of moving the MegaPrime VME crate up to the prime focus control unit GFIC rack. (Will) Fabrication of Environmental Controller continues. Mechanicals completed. Now in process of doing all of the wiring. (Grant) 2.3.4 TCS (III & IV) (Bill C) Engineering of new version 1_6. System worked well in almost all respects. Developed guider frame averaging during engineering night. Seems to reduce oscillations at PF. Corrected several small problems introduced in new version. Committed system revisions into CVS, to bring up to date. Wrote technical manual sections on using CVS and writing new EPICS records. Continued working on cass bonnette port to TCS IV IOC. 2.3.5 Summit infrastructure: Design, draw, fabricate and install new cass crate locking system prototype at cass crate position #3. It works well and am fabricating remaining locks as time permits. (Bobby) Troubleshoot mirror cover open light not coming on. Started on Wednesday, but after glycol accident in aluminizing room I continued on Monday morning. Found intermittent wire from JBTC2 to control console. Repaired wire and tested system OK. (Bobby) 2.3.6 Instrument Setups, Exchanges, and Tests: F8/AOB/KIR to PF/12K. (Bobby, Les, Grant) EEV2: Installed EEV2 and its related hardware on to Gecko. (Les) 2.4 SOFTWARE GROUP (B. GRUNDSETH) 2.4.1 Data Acquisition: (B. Grundseth) AOB - J. Thomas Run support. MegaPrime - J. Thomas Documentation for VPDR, VPDR. DAO WFS software FDR. MPC hardware PFDR. MP computing environment scope meeting. TCS-IV - J. Thomas Various design discussions with Bill. TCS agent discussions. Miscellaneous: Attended TOM, Gino interview. (J. Thomas, B. Grundseth) HP backups. (J. Thomas) SYSADMIN - S. Isani Created account for Simon Kras. Helped Kanoa with debugging the modem pool. Installed same version of Bash as on HP/Sun on the Linux hosts. Installed CDROM and sound drivers on the Linux hosts. "Up"graded to 128-bit strong encryption netscape 4.72. Ordered Dell 420's and PC parts from buy.com. NEO - S. Isani Design work on status server layout. Design work on new FITS header library, libfh. Discussed using status server as FITS header cache with Jim. Updated some EPICS variable names in the Web-TCS status screen. Added 1Hz refresh view of the guider box (animated GIF) to the Web-TCS window. See http://neptune:911/ (at night when we have a guide star.) Caution: your netscape may bloat and crash! NEO - R. Alles Continuing work on the TCSIV side for the TCS agent. Continuing work on the UNIX side for the TCS agent. QSO - Renaud Savalle Memo "Communication scheme between QSO and NEO": discussed scripts calls and real-time issues with SI and finished document. PH2 (Phase 2 Tool) - Renaud Savalle Proofread memo QSO-007 (PH2 Specs and Design) by PM. Investigated JavaScript issues, Java Applet solution for FOV tool, made proto applet. Investigated DSS viewers (ESO Skycat, CDS Aladin) and GSC browser (gsc_browser at CFHT). Worked on PH2DB Conceptual Data Model. Linux ASE (Sybase) - Renaud Savalle Called Sybase Tech Support about pb registering for Case Xpress on the web (Case # 10622005). Archive - Renaud Savalle Wrote archive daemon dbd to archive FITS headers of CFH12K exposures into Sybase database "operations". Installed daemon into the archive pipeline. Used db and Excel to produce chart and histograms for current run. Meetings - Renaud Savalle Attended Soft Group meeting 04/20 - Soft development rules. Attended PH2 Design Review 05/02 with Astros. Safety - Renaud Savalle Attended CPR/AED/first aid course. 2.4.2 Data Reduction Facility: (R. Link) Helped set up machine, kapu, to be a tape copying, archive tape testing, and archive machine backup machine; it has 4 different types of drives connected and a large disk buffer area. Started writing documentation of all systems administration tasks and information. Continued discussing various systems administration tasks with Kanoa. Reorganized performance gathering script on several machines. Confirmed new Sun maintenance contract with a Sun rep. Troubleshot losses of Internet access. Wrote up POs for various items; monitored PO purchases; stocked supplies. Troubleshot printer problems. Managed archive system activities. Managed runid accounts. Managed remote access, security machines. Managed CFHT WWW site. Modified network information files to reflect changes in personnel/hardware. Managed observer data reduction accounts. Performed network disk space clean-up. Performed network backups. Responded to myriad user requests/problems. 2.5 OPTICS & INSTRUMENTATION (G. BARRICK) 2.5.1 CFHT-IR: Completed aux control front panel mechanical and electronic layout. (J. Ward) Ordered front panel parts. Continued FPGA fit work. (J. Ward) 2.5.2 GECKO: Began setup for next run. Attended CAF planning meeting to prepare for delivery and testing of CAF. (J. Ward) 2.5.3 CFHT-12K: Provided electronic support for new shutter work. (W. Knight, J. Ward) 2.5.4 Mirror Cover: Provided support for mirror cover electronics to troubleshoot switches. (J. Ward) 2.5.5 AOB: Continued working on getting spare APD's. Was able to track down the original spec sheets from EG&G to compare characteristics with the new model. AOB to bed and the two spare APD's tested while on the mezz. Also reordered, from another supplier, the ST style fiber transmitters after the original PO was canceled by vendor. 2.5.6 Filters: Continued working on Filter Test Stand with new ideas from Greg. 2.5.7 LAMA: Completed a step-by-step troubleshooting guide for the common LAMA problems to be left at the summit. 2.5.8 Meetings: Attended the All Staff meeting and Ken's retirement party. 3.0 SCHEDULE Thur., May 04 CAFE to summit R. Wood TCS tests - telescope motion W. Cruise Fri., May 05 FSA FDR document review at 9:00 W. Rambold/ J. Thomas/ D. Sabin/ R. Calder 12k shutter design review at 3:00 W. Knight/ G. Matsushige/ R. Atapattu/ G. Barrick/ R. Calder/ B. Starr/ J. Ward MegaPrime meeting at DAO D. Salmon/ C. Veillet Sat., May 06 SAC in Victoria (through Sunday, May 7) G. Fahlman/ C. Veillet/ D. Salmon Recycle Hawaii Board Meeting 8:00-1:00 R. Alles in the large conference room Sun., May 07 SAC meeting Mon., May 08 CAFE check tests off telescope (through G. Barrick/ Tue., May 9) N. Manset/ J. Ward Tue., May 09 Audit Committee meeting in the large C. Berthoud/ conference room from 9:00-17:00 D. Warren Wed., May 10 CAFE fit check at Cass Bonnette G. Barrick/ J. Ward 12k, PF bonnette, rotator, WFC out Daycrew M2 into PF CFH12k to 2nd floor outer coude R. Song/ L. Mizuba Gecko red install / train alignment G. Barrick/ J. Ward TIC = G. Barrick SA = N. Manset Audit Committee meeting in the large C. Berthoud/ conference room from 9:00-10:45 D. Warren Seminar by H. Dole on Cosmology with ISO F. Menard at 11:00 Audit Committee summit visit in the C. Berthoud/ afternoon D. Warren Thur., May 11 Audit Committee meeting in the large C. Berthoud/ conference room from 9:00-17:00 D. Warren FSA FDR in the Waimea Observing room J. Thomas/ at 9:00 W. Rambold/ R. Calder/ G. Barrick/ C. Veillet/ D. Salmon/ D. Sabin Fri., May 12 Grif video conference at 8:00 G. Fahlman/ O. Lai/ R. Calder/ B. Starr/ D. Salmon/ B. Grundseth/ G. Barrick Espadons document review at 10:30 G. Fahlman/ D. Salmon/ B. Grundseth/ R. Atapattu/ C. Veillet/ G. Barrick/ R. Calder/ W. Cruise/ N. Manset R. Link farewell party at 15:00 Mon., May 15 Group & Project Heads meeting at 9:00 D. Salmon/ et al. TOM at 10:30 Mega Cable PDR document in-house review W. Rambold/ at 14:00 G. Matsushige/ R. Atapattu/ D. Sabin/ D. Salmon/ R. Calder/ C. Veillet/ W. Cruise/ G. Barrick Tue., May 16 Seminar by J. Matthews at 10:30 Wed., May 17 Mega cable PDR at 9:00 W. Rambold/ G. Matsushige/ R. Atapattu/ D. Sabin/ D. Salmon/ R. Calder/ C. Veillet/ W. Cruise/ G. Barrick Open House planning meeting at 15:00 D. Salmon/ et al. 12k shutter (old) to Waimea R. Song/ L. Mizuba 12k shutter tests in Waimea W. Knight 12k prep L. Mizuba/ R. Song Thur., May 18 Espadons kick-off meeting at 9:00 (through G. Fahlman/ Friday, May 19) N. Manset/ C. Veillet/ G. Barrick/ R. Calder/ W. Cruise/ R. Atapattu/ D. Salmon/ B. Grundseth 12k prep continues R. Song/ L. Mizuba Fri., May 19 Espadons meeting continues at 9:00 Espadons summit visit in the afternoon Mon., May 22 CPR/First Aid course from 8:00-5:00 R. Atapattu/ in the Large Conference Room N. Manset/ G. Barrick/ J. Ward/ D. Salmon/ R. Calder/ E. Magnier Tue., May 23 Coude off; PF on Daycrew M2 out; WFC, PFB, rotator in Daycrew CFH12 install Daycrew/ L. Mizuba/ R. Song/ B. Starr/ J. Ward 12k filter clean J. Ward Coude train cover J. Ward EEV-2 @ coude stays cold TIC = B. Starr SA = C. Veillet CAFE interlocks (through Wed., May 24) G. Matsushige/ J. Ward Wed., May 24 KIR connect to Dione Dione crash test (1 week effort) J.C. Cuillandre/ J. Thomas Thur., May 25 Open House planning meeting at 10:00 D. Salmon/ et al. Fri., May 26 CASCA meeting Mon., May 29 Memorial Day Holiday Tue., May 30 Group & Project Heads Meeting TOM at 10:30 First Responder refresher course at the R. Atapattu/ summit at 11:00 a.m. Daycrew/ O.A.'s CAFE session software test (through Thur., G. Barrick/ June 1) B. Grundseth Wed., May 31 FTS prep & scan R. Calder/ B. Grundseth Redeye prep - no filter change R. Song/ L. Mizuba/ R. Calder PFUE model delivery to Waimea shop Thur., Jun 1 Visit from Scott Roberts from NRC for C. Veillet/ MegaPrime GFSU mechanical detail review D. Salmon/ And CCD preamp location J. Thomas/ W. Rambold/ R. Calder/ G. Barrick/ Fri., Jun 2 CAFE preparation G. Barrick/ J. Ward/ N. Manset Mon., Jun 5 KIR filter list neded O. Lai James Bixler's first day with the Optics Group Wed., Jun 7 KIR prep R. Song/ L. Mizuba Thur., Jun 8 PF off; F35 on Daycrew Cass bonnette optics change J. Ward/ J. Bixler CFH12k to bed R. Song/ L. Mizuba FTS BEAR install R. Calder/ B. Grundseth/ R. Song/ Daycrew Open House planning meeting at 10:00 D. Salmon/ et al. Gecko setup (through Fri., June 9) G. Barrick/ J. Ward TIC = R. Calder SA = T. Forveille Fri., Jun 9 KIR onto AOB R. Song/ L. Mizuba Mon., Jun 12 King Kamehameha holiday FTS InGaAs install R. Calder/ B. Grundseth/ R. Song Tue., Jun 13 Group & Project Heads meeting at 9:00 TOM at 10:30 Gecko setup (through Thur., June 15) G. Barrick/ J. Ward AOB setup (through Wed., June 14) G. Barrick/ J. Thomas Wed., Jun 14 Mega-computer/software status review J. Thomas/ B. Grundseth/ S. Isani/ C. Veillet/ D. Salmon/ R. Calder/ W. Rambold/ W. Cruise 12k prep L. Mizuba/ R. Song Thur., Jun 15 Mauna Kea Support Services Oversight D. Salmon/ Committee Meeting at 10:00 at HP S. Wood Fri., Jun 16 F35 off; F8 on Daycrew Cass bonnette optics J. Ward/ J. Bixler FTS + InGaAs off Daycrew/ R. Song AOB IR install Daycrew/ L.Mizuba/ G. Matsushige/ J. Ward/ J. Thomas/ J. Bixler/ G. Barrick TIC = J. Thomas SA = O. Lai Mon., Jun 19 CAFE align to telescope, flexure tests G. Barrick/ and daytime observing (through Tue., G. Matsushige/ June 20) J. Ward/ N. Manset Tue., Jun 20 Filter list for KIR needed F. Menard Wed., Jun 21 Engineering night: « night CAFE tests G. Barrick/ N. Manset/ J. Ward « night AOB testing J. Thomas Thur., Jun 22 Open House planning meeting at 10:00 D. Salmon/ et al. Engineering night: « night CAFE tests G. Barrick/ N. Manset/ J. Ward « night TCSIV/AOB J. Thomas Fri., Jun 23 F8 off; PF on Daycrew AOB to bed T. Szarlan CFHT12k install Daycrew/ L. Mizuba/ R. Song/ J.C. Cuillandre/ KIR left on AOB - cold? R. Song/ L. Mizuba 12k filter clean T. Szarlan/ J. Bixler TIC = J.C. Cuillandre SA = G. Fahlman Discretionary night Mon., Jun 26 Group & Project Heads meeting at 9:00 TOM at 10:30 Tue., Jun 27 KIR prep L. Mizuba/ R. Song Wed., Jun 28 KIR onto AOB L. Mizuba/ R. Song Thur., Jun 29 AOB setup T. Szarlan/ J. Thomas 2-month scheduling review at 15:00 Fri., Jun 30 AOB setup T. Szarlan/ J. Thomas Sat., Jul 1 Canada Day! Tue., Jul 4 Independence Day Holiday!! Wed., Jul 5 PF off; F8 on Daycrew CFH12k to bed L. Mizuba/ R. Song AOB IR install J. Thomas/ G. Barrick/ T. Szarlan/ G. Matsushige/ Daycrew TIC = S. Isani SA = F. Menard Thur., Jul 6 Mega cable FDR at 10:00 W. Rambold/ G. Matsushige/ R. Atapattu/ D. Sabin/ D. Salmon/ R. Calder/ C. Veillet Mon., Jul 10 Group & Project Heads meeting at 9:00 TOM at 10:30 Discretionary night Tue., Jul 11 « night discretionary « night engineering: AOB/TCS Wed., Jul 12 Engineering: TCS Thur., Jul 13 Open House planning meeting at 10:00 D. Salmon/ et al. « night discretionary « night engineering: TCS 4.0 E M E R G E N C Y S T A N D B Y S C H E D U L E ( (C)ellular phone numbers & (P)ager numbers ) Tom Beck 1-10 May Ralph Taroma 11-20 May Daycrew 898-1449 (P) 969-1201 (Home) 898-5828 (P) 959-4476 (Home) Barry Starr 1-31 May Detectors 899-5007 (Page) 883-3402 (Home) Bob Calder 1-31 May Electronics 899-8041 (Page) 776-1324 (Home) Rohendra Atapattu Operations 936-0579 (C) 775-9122 (Home) Greg Barrick 1-31 May Optics 925-7910 (P) 883-9681 (Home) Rohendra Atapattu Safety 936-0579 (C) Sidik Isani 21 April - 9 May Bernt Grundseth 10-22 May Software 925-7918 (P) 883-2307 (Home) 925-7916 (P) 887-2800 (Home) Eugene Magnier 24 April - 2 May Greg Fahlman 3-9 May Nadine Manset 10-22 May Support Astronomer 989-3603 (C) 925-7473 (P) 936-0885 (C) 899-8565 (P) 936-0885 (C) 925-7473 (P) David Woodworth 1-6 May John McDonald 5-11 May David Woodworth 11-15 May Observing Assistant 961-2630 (Summit) "" "" 5.0 C F H T C A L E N D A R COMINGS AND GOINGS J.C. CUILLANDRE - MTGS/VAC APR 27 - MAY 16 FRANCE/ENGLAND D. SALMON - SAC MEETING MAY 03 - MAY 10 CANADA G. FAHLMAN - SAC MEETING MAY 04 - MAY 10 CANADA C. VEILLET - SAC MEETING MAY 04 - MAY 08 CANADA D. WARREN - VACATION MAY 05 L. FISCHER - VACATION MAY 12 - MAY 19 D. WARREN - MEETING MAY 12 - MAY 17 COLORADO C. BERTHOUD - MEETING MAY 15 - MAY 18 COLORADO D. WOODWORTH - VACATION MAY 15 - MAY 20 D. WARREN - VACATION MAY 23 - JUN 09 W. CRUISE - VACATION MAY 30 - JUL 03 T. SZARLAN - VACATION JUN 05 - JUN 20 C. HICKMAN - VACATION JUN 07 - JUN 30 W. PEXA - VACATION JUN 09 - JUL 07 J. ERIKSEN - VACATION JUL 03 E. BRYSON - BUS/VAC JUL 07 - JUL 18 G. MATSUSHIGE - VACATION JUL 20 - AUG 04 DAS/sw/054 File: 02-02